# Dynamics of Pseudoentanglement

### v1

@@ IV.3.1 First moment
with the only difference from the case of Sec. IV.2 (with $m=1$) being the different idling time between hops (4 instead of $D/2$). Just like in Sec. IV.2, the random walk exhibits a cutoff Diaconis et al. (1990), with the approach to equilibrium taking place in a relatively narrow time window, $t\sim N[\log N\pm O(1)]$. Intuitively, to achieve equilibrium the walker has to hop in all $N$ directions with high probability (equivalently, all $N$ qubits need to be flipped with high probability). This is an instance of the so-called “coupon collector’s problem”, which gives a scaling $\sim N\log(N)$. Note that time $t$ is defined as total number of gates acting on the system; defining a rescaled time $\tau\equiv t/N$ such that each qubit flips of order once per unit $\tau$, equilibration happens at $\tau\sim\log(N)\pm O(1)$.
IV.3.2 Higher moments: Relaxation time
For higher moments $m>1$, unlike the all-to-all model of Sec. IV.2, the Markov chain on $\Sigma_{m}$ is not exactly solvable (to our knowledge). To make progress we resort to a combination of rigorous bounds and numerical calculations.
@@ IV.3.2 Higher moments: Relaxation time
However, for $m\geq 2$, we observe a new second-largest eigenvalue in the spectrum of $\Gamma$, $11-1/2N<\lambda_{1}<$ 1, giving longer relaxation times $t_{\rm rel}^{(m)}>t_{\rm rel}^{(1)}$. This is unlike the all-to-all model of Sec. IV.2, where Aldous’ spectral gap conjecture Eq. (10) (see Sec. II.3) forces $t_{\rm rel}^{(m)}$ to be $m$-independent. Nonetheless, Fig. 6(a) shows no further $m$-dependence in $\lambda_{1}$ beyond $m=2$. This suggests that $m=1$ is special in this problem.
By numerical inspection, we find that the new slowest mode of $\Gamma$ for $m=2$, the function $f_{1}(S)$ such that $\Gamma f_{1}=\lambda_{1}f_{1}$, depends on the subset $S=\{\mathbf{z},\mathbf{z}^{\prime}\}$ only through the “relative coordinate” $\mathbf{r}\equiv\mathbf{z}\oplus\mathbf{z}^{\prime}$. In Appendix D.3 we show that the subspace $R$ spanned by functions that depend on $S$ only through the relative coordinate $\mathbf{r}$ is invariant under the Markov chain. Thus we can block-diagonalize $\Gamma=\Gamma|_{R}\oplus\Gamma|_{R^{\perp}}$, where $\Gamma|_{R}$ describes a random walk of $\mathbf{r}$. Numerical diagonalization confirms that the leading eigenvalue $\lambda_{1}$ of $\Gamma$ indeed comes from the $\Gamma|_{R}$ block: results for $\lambda_{1}$ agree to numerical precision between the two Markov chains. Furthermore, due to the reduced dimensionality ($\sim D$ for $\Gamma|_{R}$ compared to $\sim D^{2}$ for the full $\Gamma$), we are able to numerically diagonalize the restricted Markov chain up to larger values of system size $N$. Results, shown in Fig. 6(b), indicate a clear linear scaling $t_{\rm rel}^{(2)}=\Theta(N)$.
@@
which we believe is very loose. Proving (or disproving) our Conjecture 1, and/or tightening bounds around the relaxation times $t_{\rm rel}^{(m)}$, are left as goals for future research. However, since a precise determination of relaxation times in the arguably simpler models of Ref. Gowers (1996); Hoory et al. (2005); Brodsky and Hoory (2008) is still lacking, we suspect these may be hard problems.
IV.3.3 Higher moments: Mixing time
@@
Assuming the validity of Conjecture 1, $t_{\rm rel}^{(m)}=\Theta(N)$, we move on to the question of mixing times. A general two-sided bound on the mixing time in terms of the relaxation time is given by Jerison (2013); Levin and Peres (2017)
@@
where $\epsilon$ is the accuracy threshold on total variation distance used to define mixing; see Sec. II.2. We take $\epsilon=1/4$ in the following. Plugging Conjecture 1 in Eq. (28) yields the asymptotic bounds
@@
The lower bound Eq. (28) clearly cannot be used to gap $t_{\rm mix}^{(m)}$ away from $t_{\rm rel}^{(m)}$, which is the necessary condition for a cut-off phenomenon. To study the possible presence of a cut-off we must find stronger lower bounds on the mixing time. We show two such bounds next.
Fact IV.1 (counting bound).
For all $m$, the mixing for the Markov chain induced by the local gate set $\mathcal{G}$ obeys
@@
Let us count the number of states that can be reached from an initial state $S_{0}\in\Sigma_{m}$ in $t$ time steps. As each state is connected to at most $|\mathcal{G}|=4N$ different states, the volume of a ball of radius $t$ in $\Sigma_{m}$ (viewed here as a graph with edges given by the action of $\mathcal{G}$) is at most $(4N)^{t}$. To achieve mixing, the distribution $p_{t}$ needs to be supported on, say, at least $1/2$ of the states. Thus we must have $(4N)^{t}\geq|\Sigma_{m}|/2$, giving $t\log(4N)\geq\log\binom{D}{m}\geq mN$. ∎
Fact IV.2 (locality bound).
For all $m\geq 2$, the mixing for the Markov chain induced by the local gate set $\mathcal{G}$ on a 1D qubit chain obeys
@@
It is easy to see that, starting from a subset $S_{0}$ where all bitstrings agree outside the subsystem $A$ ($z_{j}=0$ for all sites $j=1,\dots N-N_{A}$, for all bitstrings $\mathbf{z}\in S_{0}$), the region outside a “light cone” emanating from $A$ remains in a product state: local automaton gates cannot create superpositions, so coherence must spread outward from $A$ one gate at a time. As the gate sequence is random, the light cone fluctuates instance by instance, however its width is bounded above111111 We can bound $\ell(t)$ above by noting that $\ell(t+1)=\ell(t)+1$ with probabiliy at most $2/N$ (the $\textsf{CC}X$ gate must have the target qubit $i$ at either light cone front), and $\ell(t+1)\leq\ell(t)$ otherwise. Then $\overline{\delta\ell}\leq 2/N$ and thus $\ell(t)\leq\ell(0)+2t/N$ with finite probability. This gives an upper bound on the butterfly velocity $v\leq 1$. by $\ell(t)\leq N_{A}+2vt/N$, for some constant ‘butterfly velocity’ $v$, with finite probability. Whenever $\ell(t)<N$, the distribution $p_{t}(S)$ is supported only on a small fraction of the state space $\Sigma_{m}$ (all subsets $S$ that correspond to a product state outside the light cone), and thus must be far from uniform121212 Another way to argue this is by the distinguishing statistics method in Ref. Levin and Peres (2017) applied to, e.g., $\overline{\langle Z_{i}\rangle^{2}}$ for a qubit $i$ far from $A$. . Equilibrium can only be reached after the light cone has spread across the whole system, giving $\ell(t)\geq\Omega(N)$ or $t\geq\Omega(N^{2})$. ∎
We note that, while the counting bound (Eq 30) applies to all $m$, the locality bound Eq. 31) does not apply to $m=1$ (indeed we have $t_{\rm mix}^{(1)}\sim N\log N$, Eq. (26)). This is because “subsets” with $m=1$ are just individual bitstrings, so equilibration does not require coherence (i.e. superpositions) to spread across the system. This physical difference in the effects of locality is illustrated in Fig. 7. We show results of numerical simulations for circuit-averaged quantities $\overline{\langle Z_{i}\rangle}$ (a first-moment quantity that can equilibrate under local bitflips) in Fig. 7(a) and $\overline{\langle Z_{i}\rangle^{2}}$ (a second-moment quantity that requires superpositions to equilibrate) in Fig. 7(b). Again we recall that “time” $t$ here is measured as total number of gates in the circuit, so $t\sim N^{2}$ here denotes ballistic, not diffusive, spreading of correlations.
Putting Facts IV.1 and IV.2 together with the upper bound Eq. (29), we arrive at
@@
These bounds are summarized in Fig. 5(b). We remark again that the upper bound Eq. (29) assumes Conjecture 1; an unconditional, likely loose bound $t_{\rm mix}^{(m)}\leq O(m^{2}N^{6})$ can be derived based on Eq. (27) (see Appendix F).
@@
[…]
Figure 8: Exact numerical simulation of the Markov chain induced by the local automaton gate set on $N=5$ qubits for subset sizes $m=1,\dots 8$. The initial quantum state is taken to be |$\ket{00+++}$, a subset state with $K=8$. The total variation distance from uniformity decays as $\|p_{t}-\pi_{m}\|_{\rm tv}\sim\lambda^{t-\delta t}$ at late times (bottom panel, dashed lines). Fit results for $\lambda$ (not shown) are consistent with the results of Fig. 6, with $\lambda\simeq 0.90$ for $m=1$ and $\lambda\simeq 0.92$ for all $m\geq 2$; the temporal offset $\delta t$, a proxy for the mixing time (up to additive constants), is found to scale linearly in $m$ (bottom panel, inset).
@@
    For constant $m$, we have $t_{\rm mix}^{(m)}=\Theta(N^{2})$, and thus the pseudothermalization time
    $t_{\rm PTherm}^{(m)}\sim N^{2}[1\pm O(1/N)],$ (34)
    with the $\pm$ term denoting the width of the time window in which the transition to equilibrium happens. This suggests a picture based on a ‘pseudothermalization front’ spreading ballistically through the system, with fast equilibration behind the front, so that $m$-copy pseudothermalization happens suddenly, and simultaneously for all constant $m$, once the front has swept across the system.
@@
    In the intermediate range $\omega(1)\leq m\leq O(N\log N)$, our bounds are insufficient to pin down the scaling of $t_{\rm PTherm}^{(m)}$; we have
    $t_{\rm PTherm}^{(m)}\sim N^{2}[f(m)\pm O(1/N)],$ (35)
    for some function $\Omega(1)\leq f(m)\leq O(m)$. Depending on the scaling of $f(m)$, it is possible that the ‘pseudothermalization front’ picture ceases to apply, with a parametrically slower equilibration behind the ballistic front and a separation of consecutive pseudothermalization transitions $t_{\rm PTherm}^{(m)}$, $t_{\rm PTherm}^{(m+1)}$.
@@
    At even higher moments $\omega(N\log N)\leq m\leq O({\rm poly}(N))$, the counting bound becomes dominant over the locality bound, so the ‘font’ picture is not relevant anymore. We have
    $t_{\rm PTherm}^{(m)}\sim N[mg(N)\pm O(1)],$ (36)
    for some function $\Omega(1/\log N)\leq g(N)\leq O(N)$. Depending on the scaling of $g(N)$, consecutive pseudothermalization times $t_{\rm PTherm}^{(m)}$, $t_{\rm PTherm}^{(m+1)}$ may be sharply separated, or the two crossover windows may orverlap. These two options are sketched in Fig. 5(c).
  4. (iv)
    Finally, for $m\geq\omega({\rm poly}(N))$, we have $t_{\rm mix}^{(m)}\sim m$ up to ${\rm poly}(N)$ factors.
IV.3.4 Higher dimension
@@
For constant $d$, all the results of this section proceed unchanged, except for the locality bound, Eq. (31); there, we must replace Eq. (31) with $t_{\rm mix}^{(m)}\geq\Omega(N^{1+1/d})$, with $d$ the spatial dimension. This still meets the requirement for a cut-off, Eq. (33), but is not enough to pin down the scaling of pseudothermalization time at constant $m$: we have instead $\Omega(N^{1+1/d})\leq t_{\rm mix}^{(m)}\leq O(N^{2})$.
@@
Qualitatively, this opens the question of whether, in dimension $d\geq 2$, the ‘pseudothermalization front’ picture still holds; i.e., if the state behind the light cone equilibrates quickly (which would give $t_{\rm mix}^{(m)}\sim N^{1+1/d}$), or whether it equilibrates more slowly, in such a way that even after the light cone has swept through the whole system it is still possible to distinguish the states from Haar-random ones with a constant number of copies $m$.
V Generation of Pseudoentanglement

### v3

@@ 4.3.1 First moment
with the only difference from the case of Sec. 4.2 (with $m=1$) being the different average idling time between hops (4 instead of $D/2$). Just like in Sec. 4.2, the random walk exhibits a cutoff diaconis_asymptotic_1990 , with the approach to equilibrium taking place in a relatively narrow time window, $t\sim N[\log N\pm O(1)]$. Intuitively, to achieve equilibrium the walker has to hop in all $N$ directions with high probability (equivalently, all $N$ qubits need to be flipped with high probability). This is an instance of the so-called “coupon collector’s problem”, which gives a scaling $\sim N\log(N)$. Note that time $t$ is defined as total number of gates acting on the system; in terms of an intensive time $\tau\equiv t/N$, such that each qubit flips of order once per unit $\tau$, equilibration happens at $\tau\sim\log(N)\pm O(1)$.
4.3.2 Higher moments: Relaxation time
For higher moments $m>1$, unlike the all-to-all model of Sec. 4.2, the Markov chain on $\Sigma_{m}$ is not exactly solvable (to our knowledge). To make progress we resort to a combination of rigorous bounds and numerical calculations.
@@ 4.3.2 Higher moments: Relaxation time
However, for $m\geq 2$, we observe a new second-largest eigenvalue in the spectrum of $\Gamma$, $\lambda_{1}>1-1/2N$, giving longer relaxation times $t_{\rm rel}^{(m)}>t_{\rm rel}^{(1)}$. This is unlike the all-to-all model of Sec. 4.2, where Aldous’ spectral gap conjecture, Eq. (11) (see Sec. 2.3), forces $t_{\rm rel}^{(m)}$ to be $m$-independent. Nonetheless, Fig. 6(a) shows no further $m$-dependence in $\lambda_{1}$ beyond $m=2$.
By numerical inspection, we find that the new slowest mode of $\Gamma$ for $m=2$, the function $f_{1}(S)$ such that $\Gamma f_{1}=\lambda_{1}f_{1}$, depends on the subset $S=\{\mathbf{z},\mathbf{z}^{\prime}\}$ only through the “relative coordinate” $\mathbf{r}\equiv\mathbf{z}\oplus\mathbf{z}^{\prime}$. In Appendix D.3 we show that the subspace $R$ spanned by functions that depend on $S$ only through the relative coordinate $\mathbf{r}$ is invariant under the Markov chain. Thus we can block-diagonalize $\Gamma=\Gamma|_{R}\oplus\Gamma|_{R^{\perp}}$, where the block $\Gamma|_{R}$ describes a random walk of $\mathbf{r}$. Numerical diagonalization confirms that the leading eigenvalue $\lambda_{1}$ of $\Gamma$ comes from the $\Gamma|_{R}$ block: results for $\lambda_{1}$ agree to numerical precision between the two Markov chains. Furthermore, due to the reduced dimensionality ($\sim D$ for $\Gamma|_{R}$ compared to $\sim D^{2}$ for the full $\Gamma$), we are able to numerically diagonalize the restricted Markov chain up to larger values of system size $N$. Results, shown in Fig. 6(b), indicate a clear linear scaling $t_{\rm rel}^{(2)}=\Theta(N)$.
@@
which can be tightened to $t_{\rm rel}^{(m)}\leq\tilde{O}(mN)$ using recent results by He and O’Donnell he_pseudorandom_2024 . Yet more recent results by Chen et al. chen_incompressibility_2024 rigorously prove a result equivalent to our Conjecture 1 in models of 3-local, but not geometrically local, random reversible circuits. Adapting their result to the present setting, and thus proving our Conjecture 1, is as an interesting direction for future work.
4.3.3 Higher moments: Mixing time
@@
Assuming the validity of Conjecture 1, $t_{\rm rel}^{(m)}=\Theta(N)$, we move on to the question of mixing times. A general two-sided bound on the mixing time in terms of the relaxation time is given by jerison_general_2013 ; levin_markov_2017
@@
where $\epsilon$ is the accuracy threshold on total variation distance used to define mixing; see Sec. 2.2. We take $\epsilon=1/4$ in the following. Plugging Conjecture 1 into Eq. (29) yields the asymptotic bounds
@@
The lower bound Eq. (29) clearly cannot be used to gap $t_{\rm mix}^{(m)}$ away from $t_{\rm rel}^{(m)}$ (necessary condition for a cut-off phenomenon). Therefore, to study the possible presence of a cut-off we must find stronger lower bounds on the mixing time. We show two such bounds next.
Fact 4.1 (counting bound).
For all $m$, the mixing time of the Markov chain induced by the local gate set $\mathcal{G}$ obeys
@@
Let us count the number of states that can be reached from an initial state $S_{0}\in\Sigma_{m}$ in $t$ time steps. As each state is connected to at most $|\mathcal{G}|=4N$ different states, the volume of a ball of radius $t$ in $\Sigma_{m}$ (viewed here as a graph with edges given by the action of $\mathcal{G}$) is at most $(4N)^{t}$. To achieve mixing, the distribution $p_{t}$ needs to be supported on, say, at least $1/2$ of the elements of $\Sigma_{m}$. Thus we must have $(4N)^{t}\geq|\Sigma_{m}|/2$, giving $t\log(4N)\geq\log\binom{D}{m}-\log 2\geq\Omega(mN)$. ∎
Fact 4.2 (locality bound).
For all $m\geq 2$, the mixing time of the Markov chain induced by the local gate set $\mathcal{G}$ on a 1D qubit chain obeys
@@
It is easy to see that, starting from a subset $S_{0}$ where all bitstrings agree outside the subsystem $A$ ($z_{j}=0$ for all sites $j=1,\dots N-N_{A}$, for all bitstrings $\mathbf{z}\in S_{0}$), the region outside a “light cone” emanating from $A$ remains in a computational basis state: local automaton gates cannot create superpositions from such states, so coherence must spread outward from $A$ one gate at a time. As the gate sequence is random, the light cone fluctuates instance by instance, however its width $\ell(t)$ is bounded above161616 We can bound $\ell(t)$ above by noting that $\ell(t+1)=\ell(t)+1$ with probabiliy at most $2/N$ (the $\textsf{CC}X$ gate must have the target qubit $i$ at either light cone front), and $\ell(t+1)\leq\ell(t)$ otherwise. Then $\overline{\delta\ell}\leq 2/N$ and thus $\ell(t)\leq\ell(0)+2t/N$ with finite probability. This gives an upper bound on the butterfly velocity $v\leq 1$. by $\ell(t)\leq N_{A}+2vt/N$, for some constant ‘butterfly velocity’ $v$, with finite probability. Whenever $\ell(t)<N$, the distribution $p_{t}(S)$ is supported only on a small fraction of the state space $\Sigma_{m}$ (all subsets $S$ that correspond to a product state outside the light cone), and thus must be far from uniform171717 Another way to argue this is by the distinguishing statistics method in Ref. levin_markov_2017 applied to, e.g., $\overline{\langle Z_{i}\rangle^{2}}$ for a qubit $i$ far from $A$. . Equilibrium can only be reached after the light cone has spread across the whole system, giving $\ell(t)\geq\Omega(N)$ or $t\geq\Omega(N^{2})$. ∎
We note that, while the counting bound (Eq 31) applies to all $m$, the locality bound (Eq. 32) does not apply to $m=1$; indeed we have $t_{\rm mix}^{(1)}\sim N\log N$, Eq. (27). This is because “subsets” with $m=1$ are just individual bitstrings, so equilibration does not require coherence (i.e. superpositions) to spread across the system. This physical difference in the effects of locality is illustrated in Fig. 7. We show results of numerical simulations for circuit-averaged quantities $\overline{\langle Z_{i}\rangle}$ (a first-moment quantity that can equilibrate under local bitflips) in Fig. 7(a) and $\overline{\langle Z_{i}\rangle^{2}}$ (a second-moment quantity that requires superpositions to equilibrate) in Fig. 7(b). The overline denotes averaging over instances of the random circuit. “Time” $t$ here is measured as total number of gates in the circuit, so $t\sim N^{2}$ here denotes ballistic, not diffusive, spreading of correlations.
Putting Facts 4.1 and 4.2 together with the upper bound Eq. (30), we arrive at181818 Note that the mixing time $t_{\rm mix}^{(m)}$ is conventionally based on a worst-case pure initial state, $p_{0}(S)=\delta_{S,S_{0}}$; in the present context, the initial state under consideration is already partially mixed, $p_{0}(S)={\rm const.}>0$ if $S$ is made of bitstrings that are identically 0 outside subsystem $A$, $p_{0}(S)=0$ otherwise (describing the initial random subset-phase ensemble on $A$, see also Appendix B). The upper bound $t_{\rm indist}^{(m)}\leq O(mN^{2})$ holds a fortiori for our mixed initial state; the locality lower bound $t_{\rm indist}^{(m)}\geq\Omega(N^{2})$ and the counting lower bound $t_{\rm indist}^{(m)}\geq\Omega(mN/\log N)$ also apply without changes as long as $N_{A}=o(N)$.
@@
These bounds are summarized in Fig. 5(b). We remark again that the upper bound Eq. (30) assumes Conjecture 1; an unconditional bound $t_{\rm mix}^{(m)}\leq O(m^{2}N^{6})$ can be derived based on Eq. (28) (see Appendix F) and can be further tightened to $t_{\rm mix}^{(m)}\leq\tilde{O}(m^{2}N^{2})$ using recent results he_pseudorandom_2024 .
@@
which is a necessary condition for a cut-off, see Sec. 2.2. We thus conjecture the presence of asymptotically sharp distinguishability transitions in this model. While numerical simulations are strongly limited by the explosive growth of the state space $|\Sigma_{m}|=\binom{2^{N}}{m}$, results of exact simulations of $N=5$ qubits, shown in Fig. 8, are consistent with the emergence of a cutoff—we clearly see the development of a plateau, followed by a decay $\sim\lambda_{1}^{t-\delta t}$ of the total variation distance, with a temporal offset $\delta t\sim m$ in the accessible range $1\leq m\leq 8$ (for $N=5$, $m=8$ we have $|\Sigma_{8}|=\binom{32}{8}\gtrsim 10^{7}$ states).
Figure 8: Exact numerical simulation of the Markov chain induced by the local automaton gate set (Sec. 4.3) on $N=5$ qubits for subset sizes $m=1,\dots 8$. The initial quantum state is taken to be |$\ket{00+++}$, a subset state with $K=8$. The total variation distance from uniformity decays as $D_{\rm tv}(p_{t},\pi_{m})\sim\lambda^{t-\delta t}$ at late times (bottom panel, dashed lines). Fit results for $\lambda$ (not shown) are consistent with the results of Fig. 6, with $\lambda\simeq 0.90$ for $m=1$ and $\lambda\simeq 0.92$ for all $m\geq 2$; the temporal offset $\delta t$, a proxy for the mixing time (up to additive constants), is found to scale linearly in $m$ (bottom panel, inset).
@@
    For constant $m$, we have $t_{\rm mix}^{(m)}=\Theta(N^{2})$, and thus the indistinguishability time
    $t_{\rm indist}^{(m)}\sim N^{2}[1\pm O(1/N)],$ (35)
    with the $\pm$ term denoting the width of the time window in which the transition to equilibrium happens. This suggests a picture based on a ‘pseudoentanglement front’ spreading ballistically through the system, with fast equilibration behind the front. In this picture, $m$-copy indistinguishability arises suddenly, and simultaneously for all constant $m$, once the front has swept across the system.
@@
    In the intermediate range $\omega(1)\leq m\leq\tilde{O}(N)$, our bounds are insufficient to pin down the scaling of $t_{\rm indist}^{(m)}$; we have
    $t_{\rm indist}^{(m)}\sim N^{2}[f(m)\pm O(1/N)],$ (36)
    for some function $\Omega(1)\leq f(m)\leq O(m)$. Depending on the scaling of $f(m)$, it is possible that the ‘pseudoentanglement front’ picture ceases to apply, with a parametrically slower equilibration behind the ballistic front and a separation of consecutive indistinguishability transitions $t_{\rm indist}^{(m)}$, $t_{\rm indist}^{(m+1)}$.
@@
    At even higher moments $\tilde{\omega}(N)\leq m\leq O({\rm poly}(N))$ (super-linear, but still polynomial in $N$), the counting bound becomes dominant over the locality bound, so the ‘front’ picture is not relevant anymore. We have
    $t_{\rm indist}^{(m)}\sim N[mg(N)\pm O(1)],$ (37)
    for some function $\tilde{\Omega}(1)\leq g(N)\leq O(N)$. Depending on the scaling of $g(N)$, consecutive indistinguishability times $t_{\rm indist}^{(m)}$, $t_{\rm indist}^{(m+1)}$ may be sharply separated, or the two crossover windows may overlap. These two options are sketched in Fig. 5(c).
4.3.4 Higher dimension
@@
For constant $d$, all the results of this section proceed unchanged, except for the locality bound, Eq. (32); there, we must replace Eq. (32) with $t_{\rm mix}^{(m)}\geq\Omega(N^{1+1/d})$, with $d$ the spatial dimension. This still meets the requirement for a cut-off, Eq. (34), but is not enough to pin down the scaling of indistinguishability time at constant $m$: we have instead $\Omega(N^{1+1/d})\leq t_{\rm mix}^{(m)}\leq O(N^{2})$.
@@
Qualitatively, this opens the question of whether, in dimension $d\geq 2$, the ‘pseudoentanglement front’ picture still holds; i.e., if the state behind the light cone equilibrates quickly (which would give $t_{\rm mix}^{(m)}\sim N^{1+1/d}$), or whether it equilibrates more slowly, in such a way that even after the light cone has swept through the whole system it is still possible to distinguish the states from Haar-random ones with a constant number of copies $m$. We leave this question for future work.
5 Generation of Pseudoentanglement